# A Quantum Approach to Synthetic Minority Oversampling Technique (SMOTE)

### v2

@@ IV-A7 Final thoughts on SMOTE Performance
The Random Forest model demonstrates good performance, regardless of SMOTE. This underscores the model’s intrinsic abilities in effectively dealing with class imbalances [18]. The ensemble strategy of the model, which combines predictions from numerous decision trees, inevitably offers a degree of resilience to imbalance, which is further strengthened by the use of SMOTE. Fernandez et al. [5] provide evidence supporting the effectiveness of ensemble approaches in handling unbalanced data. They propose that combining techniques such as Random Forest with SMOTE may lead to substantial improvements in model performance. All of the findings described in the assessment of Model performances are summarized in the table I and the Confusion Matrix comparison table II.
IV-B Inferences from Simulation
@@ Acknowledgment
The authors are grateful to the IBM Quantum Experience platform and their team for developing the Qiskit platform and providing open access to their simulators for running quantum circuits and performing the experiments reported here. Authors also express gratitude towards Center for Quantum Software and Information (CQSI) and Sydney Quantum Academy.

### v3

@@ 4.1.7 Final thoughts on SMOTE Performance
The Random Forest model demonstrates good performance, regardless of SMOTE. This underscores the model’s intrinsic abilities in effectively dealing with class imbalances [19]. The ensemble strategy of the model, which combines predictions from numerous decision trees, inevitably offers a degree of resilience to imbalance, which is further strengthened by the use of SMOTE. Fernandez et al. [5] provide evidence supporting the effectiveness of ensemble approaches in handling unbalanced data. They propose that combining techniques such as Random Forest with SMOTE may lead to substantial improvements in model performance. All of the findings described in the assessment of Model performances are summarized in the table 1 and the Confusion Matrix comparison table 2.
4.2 Inferences from Simulation
@@ Acknowledgment
The authors are grateful to the IBM Quantum Experience platform and their team for developing the Qiskit platform and providing open access to their simulators for running quantum circuits and performing the experiments reported here. The authors also express gratitude towards the Center for Quantum Software and Information (CQSI) and Sydney Quantum Academy.
6 Statements and Declarations
Competing Interests: The authors have no financial or non-financial competing interests.
Authors’ contributions: The authors confirm their contribution to the paper as follows: Study conception and design: N.M., B.K.B., C.F., P.D.;
Data collection: N.M.;
Analysis and interpretation of results: N.M., B.K.B., C.F., P.D.;
Draft manuscript preparation: N.M., B.K.B., C.F., P.D.;
All authors reviewed the results and approved the final version of the manuscript.
Funding: Authors declare that there has been no external funding.
Availability of data and materials: All the data provided in this manuscript is generated during the simulation and can be provided upon reasonable request.